The sidecap package (text in body of floats)

If you need to put text inside the body of a floating environment, you’ll have to put it inside a \parbox or a minipage environment. (This is necessary due to the way the sidecap package grabs the body of the float.)

For example, the first pair of floats in this document (fig. 1 and tab. 1) contain essentially the following:

\begin{minipage}{0.5\textwidth}%
text text ... text text.
\end{minipage}%

Note that the alignment of the caption with respect to text in the body is wrong.

text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text.

Fig. 1: ‘SCfigure’ The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text.

Tab. 1: ‘SCtable’ The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

To fix this problem, you’ll have to use some extra effort.

For example, the second pair of floats (fig. 2 and tab. 2) contain essentially the following:

\begin{minipage}{0.5\textwidth}%
\strut
text text ... text text.
\unskip\strut
\end{minipage}%

Now the caption is properly aligned with the first or last line of the text in the float’s body.

text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text.

Fig. 2: ‘SCfigure’ The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text text.

Tab. 2: ‘SCtable’ The dazed brown fox quickly gave 12345– 67890 jumps! The dazed brown fox quickly gave 12345–67890 jumps!

Still, there is a flaw in the captions: Due to the small \linewidth the linebreaking and/or word spacing may be awful. Use one of the raggedright, raggedleft, or ragged package options.